I am just starting 7 grade please show me how to find the missing numbers n,8,18,3,6,and 3 mean =9
Answer by josmiceli(19441)

n,8,18,3,6,and 3 mean =9
(n + 8 + 18 + 3 + 6 + 3) / 6 = 9
(n + 38) / 6 = 9
n + 38 = 6*9
n = 54 - 38
n = 16